The first tern of an A.P. is 2 and the 6th term is 12. Find the

sum of the first 12 terms.


1
Expert's answer
2021-09-15T11:39:00-0400

For the 2nd and 6th terms, we have:


"a_2=a_1+d(2-1)=a_1+d,\\
a_6=a_1+d(6-1)=a_1+5d;\\
d=2.5,\\
a_1=-0.5."

The sum:


"\Sigma=\frac n2(2a_1+d(n-1))=\\\space\\
=\frac{12}{2}(2·(-0.5)+2.5(12-1))=159."